OpenStudy (anonymous):

18. Skip designs tracks for amusement park rides. For a new design, the track will be elliptical. If the ellipse is placed on a large coordinate grid with its center at (0, 0), the equation x^2/2500+y^2/8100=1 the path of the track. The units are given in yards. How long is the major axis of the track? Explain how you found the distance. (3 points)

OpenStudy (anonymous):

@terenzreignz can u help me???

terenzreignz (terenzreignz):

Why is there a (3 points) at the end? Is this a quiz? The feds aren't going to be pleased :P Nonetheless, the ellipse has the equation\[\huge \frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] So, just find a and b, and whichever is bigger, that's the length of HALF of the major axis \[\huge \frac{x^2}{2500}+\frac{y^2}{8100}=1\] Please try it now.